# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ VII Destruction of the swap phase in 2D by spiral defects
In small 2D systems, simulations show states in which $M_{A}$ and $M_{B}$ oscillate in time (like in the MF swap phase). These oscillations are irregular, and typically occur through nucleation of droplets of opposite magnetization by the “unsatisfied” species, in alternating order (see Movie 1).
@@
Despite having 4-fold rather than full rotational symmetric shape, the spiral waves observed in the NR Ising model appear to behave similarly to those emerging in the CGL equation [84, 156, 157, 98, 101, 102, 158, 159, 160, 161, 162]. Defects are created and annihilated in pairs and opposite topological charges ($+$ and -- defects) lead to opposite angular velocities (see Movie 1). Finally we note that in mean-field, a homogeneously oscillating system is stable, but given sufficiently strong added noise spirals emerge, as in our MC simulations (see Appendix C).
Quantitatively, the absence of a swap phase in the thermodynamic limit can be seen from the behavior of the order parameters $R$ and $\mathcal{L}$ (Eq. (47) and (VI), respectively) as system size increases. In Fig. 9 color maps of $R$ and $\mathcal{L}$ are shown as a function of $\tilde{J}$ and $\tilde{K}$ for both $L=20$ and $L=80$. For $L=20$, there is a region in which both $R$ and $\mathcal{L}$ are non-zero, compatible with the swap phase. However, for $L=80$, both $R$ and $\mathcal{L}$ diminish in this region, indicating instead a disorder phase. Figure 10 shows the order parameters as function of $\tilde{J}$ for fixed $\tilde{K}$, at different system sizes. The intermediate region ($1.5\lesssim\tilde{J}\lesssim 3$) corresponds to the finite-system swap state, but both $R$ and $\mathcal{L}$ approach zero as $L$ increases. In addition, the putative critical $\tilde{J}$ marking the transition from disorder to swap depends on the system size, signifying the absence of a well-defined phase transition in the thermodynamic limit.
To conclude, our numerical simulations indicate that the MF swap phase is destroyed by spiral defects in 2D where it is replaced by a disordered state.
Fig. 9: Instability of the swap phase in 2D shown by the order parameters. Color map of the (a) synchronization order parameter $R$ and (b) phase-space angular momentum $\mathcal{L}$, as a function of $\tilde{J}$ and $\tilde{K}$, for a 2D system with $L=20$. (c-d) The same with $L=80$. Obtained from MC simulations.
Fig. 10: No phase transition from disorder to swap in 2D. (a) $R$ and (b) $\mathcal{L}$, shown as a function of $\tilde{J}$ for $\tilde{K}=0.3$, for different linear system size $L$. Dashed red line shows upper bound for $R$ and $\mathcal{L}$ for points that did not converge during simulation running time. Both $R$ and $\mathcal{L}$ approach zero as $L$ increases.
@@ VIII Existence of a stable swap phase in 3D
Fig. 11: Time evolution of a stable swap phase in 3D. (a) Kymograph of $\theta$ for a single 1D row in the 3D system. (b) $M_{A}$, $M_{B}$ and $s=\sqrt{\left(M_{A}^{2}+M_{B}^{2}\right)/2}$ as a function of time. Obtained from MC simulation with parameters: $\tilde{J}=1.5$, $\tilde{K}=0.1$ and $L=160$.
In 3D, numerical simulations suggest that a stable swap phase does exist. Its behavior is demonstrated in Fig. 11 where we show a kymograph of the $\theta$-field of a single 1D row in the 3D lattice as well as plot the total magnetizations and $s$ (Eq. (47)) as a function of time. A full visualization of this phase is shown in Movie 2. Unlike the 2D droplet-induced swap persisting at finite systems (Fig. 7), here the oscillations are noisy and spatially homogeneous, without any apparent structures. Moreover, $M_{A}(t)$ and $M_{B}(t)$ display coherent oscillations with a fixed period and phase shift between the species.
Simulations with varying system sizes indicate there is a well-defined phase transition between a disordered phase with $R=\mathcal{L}=0$ to a swap phase with non-zero $R$ and $\mathcal{L}$. Figure 12 shows a color map of $R$ in the $(\tilde{K},\tilde{J})$ space, for $L=320$, while Fig. 13 shows a cut of the color map for $\tilde{K}=0.3$ with different $L$ values. Unlike in 2D (Fig. 10), here there is a critical $\tilde{J}$, below which $R$ approaches zero with $R\sim L^{-d/2}$ scaling and above which it approaches a finite non-zero value as $L$ increases, see inset of Fig. 13a. Moreover, the phase transition appears to be second-order in the sense that $R$ and $\mathcal{L}$ change continuously, and only their derivative is discontinuous.
@@
The second-order phase transition is limited to small $\tilde{J}$ and $\tilde{K}$ values. This is seen in Fig. 12 where the change in $R$ and $\mathcal{L}$ at the phase transition appears to be smooth for small ($\tilde{J}$, $\tilde{K}$) values and discontinuous for higher values. In the next section, we focus on the second-order phase transition and discuss the other type in Sec. VIII.3.
Fig. 12: Order parameters in a large 3D system. Color map of (a) $R$ and (b) $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}$ for a 3D system with $L=320$. In (a), red dot marks the pitchfork bifurcation, arrows corresponds to fixed $\tilde{K}$ and fixed $\tilde{J}$ cuts analyzed in Figs. 131416 and 20 and stars depict points analyzed in Fig. 15. Obtained from MC simulations.
Fig. 13: Phase transition from disorder to swap in 3D. (a) $R$ as a function of $\tilde{J}$ for $\tilde{K}=0.3$ shown for different linear system size $L$. Inset: $R$ vs. $L^{-1}$ for $\tilde{K}=0.3$ and $\tilde{J}=1.38,1.40,1.42,1.43,1.45$ from bottom to top on a log-log scale. (b) $\mathcal{L}$ as a function of $\tilde{J}$ with the same parameters as panel a. Obtained from MC simulations.
Fig. 14: Phase transition in 3D as a function of $\tilde{K}$. $R$ vs. $\tilde{K}$ for $\tilde{J}=1.5$ in a 3D system. Phase-space angular momentum is shown in the inset. At $\tilde{K}\approx 0.5$, a phase transition from swap to disorder occurs. Obtained from MC simulations.
@@ VIII.1 Critical exponents
Having numerically confirmed the existence of a phase transition from disorder to swap in 3D, we move on to determine the associated critical exponents of $R$. Note that $\mathcal{L}$ is not expected to yield novel critical exponents because it is proportional to $R^{2}/T_{\rm osc}$ where $T_{\rm osc}$, the oscillation period, is finite at the Hopf bifurcation (see Sec. V.2.2).
As predicted by the MF analysis, the phase-space trajectories become circular when approaching the transition line (Fig. 15), supporting a CGL-like behavior as explained in Sec. V.2.2. Moreover, renormalization group studies based on the $\epsilon$-expansion suggest that the CGL equation with additive noise has the same critical exponents as the XY model [85, 86]. Thus, in what follows we will compare our results to the 3D XY critical exponents and to the 3D Ising critical exponents corresponding to the $\tilde{K}=0$ case.
@@
According to the theory of finite-size scaling, in a finite system the order parameter $R$ and the susceptibility $\chi$, defined as $\chi=L^{d}\left(\langle s^{2}\rangle-\langle s\rangle^{2}\right)$, satisfy the following relations in the vicinity of the critical point $\tilde{J}_{c}$ (assuming fixed $\tilde{K}$[163],
$\begin{split}R&=L^{-\beta/\nu}\hat{R}(L^{1/\nu}(\tilde{J}-\tilde{J}_{c}))\\ \chi&=L^{\gamma/\nu}\hat{\chi}(L^{1/\nu}(\tilde{J}-\tilde{J}_{c})).\end{split}$ (52)
@@
Here, $\hat{\chi}(x)$ and $\hat{R}(x)$ are called the scaling functions, and $\nu$, $\gamma$ and $\beta$ are the critical exponents $\xi\sim|\tilde{J}-\tilde{J}_{c}|^{-\nu}$ where $\xi$ is the correlation length, $\chi\sim|\tilde{J}-\tilde{J}_{c}|^{-\gamma}$, and $R\sim|\tilde{J}-\tilde{J}_{c}|^{\beta}$ (for $\tilde{J}>\tilde{J}_{c}$). The scaling functions are independent of $L$ and converge to non-zero constants at the critical point, i.e. $\hat{R}(0)={\rm const_{1}}$ and $\hat{\chi}(0)={\rm const_{2}}$. In what follows we assume that the finite-size scaling theory holds out of equilibrium, as is often done [164]. We set $\tilde{K}=0.1$ where numerical evidence shows that the transition is continuous (like a second-order transition).
@@
The critical point $\tilde{J}_{c}$ is found using the Binder Cumulant [165] which is defined as $U_{L}=1-\langle s^{4}\rangle/(3\langle s^{2}\rangle^{2})$. Away from the critical point, the value of $U_{L}$ depends on system size, approaching $1/3$ in the disordered and $2/3$ in the ordered regime as $L\to\infty$ [166, 167]. However, at the critical point, in which correlation length exceeds the linear system size $L$, $U_{L}$ approaches a fixed value $U^{*}$ independent of $L$. Hence, the critical point is found from the intersection point of all the $U_{L}$ vs. $\tilde{J}$ curves plotted for different system sizes. Figure 16a shows that indeed all the curves cross at a single point ($\tilde{J}_{c}=1.34615\pm 0.00001$ for $\tilde{K}=0.1$).
@@
Next, we plot the $\chi$ and $R$ as a function of $\tilde{J}$, see Fig. 16b-c. For each system size $L$, we find the maximum $\chi_{\rm max}(L)$ of $\chi$, the value $\tilde{J}_{\rm max}(L)$ of $\tilde{J}$ at the maximum, and the value $R(\tilde{J}=\tilde{J}_{c},L)$ of $R$ at the critical point. We then use the finite-size scaling relations (which can be derived from Eq. (52[163]),
$\begin{split}\tilde{J}_{\rm max}(L)-\tilde{J}_{\rm c}&=aL^{-1/\nu}\\ \chi_{\rm max}(L)&=bL^{\gamma/\nu}\\ R(\tilde{J}=\tilde{J}_{c},L)&=cL^{-\beta/\nu}\end{split}$ (53)
@@
and perform linear fits by taking the logarithm of both equation sides (Fig. 16(d-f)). The procedure outlined here yields the critical exponents, where, to gain good accuracy, all averages are taken using $10^{6}$ independent samples of $s$. The error is estimated using the bootstrap method [163, 168]: the same procedure is repeated for multiple samples where each sample is constructed by drawing samples randomly, one at a time, from our data sample, and returning them to the data sample after they have been chosen. The standard deviation of the critical exponents is estimated by the standard deviation calculated from $10^{5}$ bootstrap resamplings of the data (Fig. 16(g-i)). Our final result is $\nu=0.68\pm 0.02$, $\gamma=1.33\pm 0.04$, and $\beta=0.35\pm 0.01$, which are in good agreement with 3D XY critical exponents $\nu_{XY}=0.67$, $\gamma_{XY}=1.32$, $\beta_{XY}=0.35$ [169, 170], more so than with 3D Ising critical exponents $\nu_{I}=0.63$, $\gamma_{I}=1.24$, $\beta_{I}=0.33$ [169], see graphical comparison in Fig. 17.
@@
Fig. 16: Calculation of the critical exponents of the 3D NR Ising model. We set $\tilde{K}=0.1$. (a-c) The Binder cumulant $U_{L}$, the susceptibility $\chi$, and the order parameter $R$ as a function of $\tilde{J}$ for linear system sizes $L=16,20,24,28,32,36,40,44,48$ (from bottom to top in panel b). The vertical dashed grey line indicates the critical point, $\tilde{J}_{c}=1.34615$. Each point is calculated by averaging over $10^{6}$ independent samples of $s$. (d-f) $\chi_{\rm max}$, $|\tilde{J}_{\rm max}-\tilde{J}_{c}|$ and $R(\tilde{J}_{c})$ as a function of $L$ on a log-log scale. Dots are extracted from analyzing panels a-c and full lines indicate best linear fits. (g-i) Standard error of $\nu$, $\gamma$, and $\beta$ as a function of the number of bootstrap samples ($10^{5}$ bootstrap resamplings are performed in total).
@@
Fig. 17: Compatibility with XY critical exponents. Critical exponents of the 3D NR Ising model (red) for $\tilde{K}=0.1$ extracted from finite-size scaling analysis along with their standard deviation represented by a semi-transparent red rectangle, and compared with the 3D Ising model (blue) and the 3D XY model (green).

### v2

@@ VII Destruction of the swap phase in 2D by spiral defects
In small 2D systems, simulations show states in which $M_{A}$ and $M_{B}$ oscillate in time (like in the mean-field swap phase). These oscillations are irregular, and typically occur through nucleation of droplets of opposite magnetization by the “unsatisfied” species, in alternating order (see Movie 1).
@@
Despite having 4-fold rather than full rotational symmetric shape, the spiral waves observed in the nonreciprocal Ising model appear to behave similarly to those emerging in the CGL equation [86, 177, 178, 102, 106, 107, 179, 180, 181, 182, 183, 184, 185]. Defects are created and annihilated in pairs and opposite topological charges ($+$ and -- defects) lead to opposite angular velocities (see Movie 1). Finally we note that in mean-field, a homogeneously oscillating system is stable, but given sufficiently strong added noise spirals emerge, as in our Monte-Carlo simulations (see Appendix D).
Quantitatively, the absence of a swap phase in the thermodynamic limit can be seen from the behavior of the order parameters $R$ and $\mathcal{L}$ (Eq. (49) and (VI), respectively) as system size increases. In Fig. 9a-d color maps of $R$ and $\mathcal{L}$ are shown as a function of $\tilde{J}$ and $\tilde{K}$ for small ($L=20$) and large ($L=80$) systems. For the $L=20$ system, there is a region in which both $R$ and $\mathcal{L}$ are non-zero, compatible with the swap phase. However, for the $L=80$ system, both $R$ and $\mathcal{L}$ diminish in this region, indicating instead a disorder phase. Figures 9e and 9f show the order parameters as function of $\tilde{J}$ for fixed $\tilde{K}$, at different system sizes. The intermediate region ($1.5\lesssim\tilde{J}\lesssim 3$) corresponds to the finite-system swap state, but both $R$ and $\mathcal{L}$ approach zero as $L$ increases. In addition, the putative critical $\tilde{J}$ marking the transition from disorder to swap depends on the system size, signifying the absence of a well-defined phase transition in the thermodynamic limit.
To conclude, our numerical simulations indicate that the mean-field swap phase is destroyed by spiral defects in 2D where it is replaced by a disordered state.
Fig. 9: Instability of the swap phase in 2D as seen by the order parameters. Color map of the (a) synchronization order parameter $R$ and (b) phase-space angular momentum $\mathcal{L}$, as a function of $\tilde{J}$ and $\tilde{K}$, for a 2D system with $L=20$. (c-d) The same with $L=80$. (e) $R$ and (f) $\mathcal{L}$, shown as a function of $\tilde{J}$ for $\tilde{K}=0.3$, for different linear system size $L$. Dashed red line shows upper bound for $R$ and $\mathcal{L}$ for points that did not converge during simulation running time. Obtained from MC simulations.
@@ VIII Existence of a stable swap phase in 3D
Fig. 10: Time evolution of a stable swap state in 3D. (a) Kymograph of $\theta$ for a single 1D row in the 3D system. (b) $M_{A}$, $M_{B}$ and $s=\sqrt{\left(M_{A}^{2}+M_{B}^{2}\right)/2}$ as a function of time. Obtained from MC simulation with parameters: $\tilde{J}=1.5$, $\tilde{K}=0.1$ and $L=160$.
In 3D, numerical simulations suggest that a stable swap phase does exist. Its behavior is demonstrated in Fig. 10 where we show a kymograph of the $\theta$-field of a single 1D row in the 3D lattice as well as a plot of the total magnetizations and $s$ (Eq. (49)) as a function of time. A full visualization of this phase is shown in Movie 2. Unlike the 2D droplet-induced swap persisting at finite systems (Movie 1), here the oscillations are noisy and spatially homogeneous, without any apparent structures. Moreover, $M_{A}(t)$ and $M_{B}(t)$ display coherent oscillations with a fixed period and phase shift between the species.
Simulations with varying system sizes indicate there is a well-defined phase transition between a disordered phase with $R=\mathcal{L}=0$ to a swap phase with non-zero $R$ and $\mathcal{L}$. Figure 11 shows color maps of $R$ and $\mathcal{L}$ in the $(\tilde{K},\tilde{J})$ space, for linear system size $L=320$, while Fig. 12a-b shows a cut of the color maps for $\tilde{K}=0.3$ with different $L$ values. The inset of Fig. 12a shows two different behaviors depending on whether $\tilde{J}$ is above or below a critical value $\tilde{J}_{c}$, unlike in 2D (Fig. 9). Below $\tilde{J}_{c}$, $R$ approaches zero with $R\sim L^{-d/2}$ scaling as $L$ increases: this corresponds to the disordered phase. Above $\tilde{J}_{c}$, $R$ converges to a finite non-zero value as $L$ increases: this corresponds to an ordered phase, here the swap phase. Moreover, the phase transition appears to be second-order in the sense that $R$ and $\mathcal{L}$ change continuously.
@@
Note that the second-order phase transition between disorder and swap is limited to small $\tilde{J}$ and $\tilde{K}$ values. At high values of $\tilde{J}$ and $\tilde{K}$, the transition appears to be discontinuous. This is discussed further in Sec. VIII.3.
Fig. 11: Order parameters in a large 3D system. Color map of (a) $R$ and (b) $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}$ for a 3D system with $L=320$. In (a), red dot marks the 3D Ising phase-transition point ($\tilde{J}_{c}\approx 1.33$ and $\tilde{K}=0$), arrows corresponds to fixed $\tilde{K}$ and fixed $\tilde{J}$ cuts analyzed in Figs. 1214 and 18 and stars depict points analyzed in Fig. 13. Obtained from MC simulations.
Fig. 12: Phase transition from disorder to swap in 3D. (a) $R$ as a function of $\tilde{J}$ for $\tilde{K}=0.3$ shown for different linear system size $L$. Inset: $R$ vs. $L^{-1}$ for $\tilde{K}=0.3$ and $\tilde{J}=1.38,1.40,1.42,1.43,1.45$ from bottom to top on a log-log scale. (b) $\mathcal{L}$ as a function of $\tilde{J}$ with the same parameters as panel a. (c) $R$ vs. $\tilde{K}$ for $\tilde{J}=1.5$ for different $L$. Phase-space angular momentum is shown in the inset. Obtained from MC simulations.
@@ VIII.1 Critical exponents
Having numerically confirmed the existence of a continuous phase transition from disorder to swap in 3D, we move on to determine the related critical exponents, focusing on the order parameter $R=\langle s\rangle$ (averages here and in the following discussion are with respect to time and realizations but for brevity, we remove the $t$ and $\Omega$ subscripts). Note that $\mathcal{L}$ is not expected to yield novel critical exponents because at the Hopf bifurcation the angular velocity in phase space is finite (see Eq. (37) for mean-field and Fig. 24 for simulation evidence), and so the phase-space angular momentum scales with $\langle s^{2}\rangle$.
As predicted by the mean-field analysis, the phase-space trajectories become circular when approaching the transition line (Fig. 13), supporting a CGL-like behavior as explained in Sec. V.2.2. Moreover, renormalization group studies based on the $\epsilon$-expansion suggest that the synchronization transition of the CGL equation with additive noise belongs to the same universality class as the phase transition from disorder to long-range order in the XY model [89, 90]. Thus, in what follows we will compare our results to the 3D XY critical exponents and to the 3D Ising critical exponents corresponding to the $\tilde{K}=0$ case.
@@
According to the theory of finite-size scaling [186, 187], in a finite system the order parameter $R$, the susceptibility $\chi$, defined as
$\chi=L^{d}\left(\langle s^{2}\rangle-\langle s\rangle^{2}\right),$ (55)
and the Binder cumulant [188] defined as
$U=1-\frac{\langle s^{4}\rangle}{3\langle s^{2}\rangle^{2}},$ (56)
satisfy the following relations in the vicinity of the critical point $\tilde{J}_{c}$ (assuming fixed $\tilde{K}$) and for large system size:
$\begin{split}R&=L^{-\beta/\nu}\hat{R}(L^{1/\nu}(\tilde{J}-\tilde{J}_{c}))\\ \chi&=L^{\gamma/\nu}\hat{\chi}(L^{1/\nu}(\tilde{J}-\tilde{J}_{c}))\\ U&=\hat{U}(L^{1/\nu}(\tilde{J}-\tilde{J}_{c})).\end{split}$ (57)
@@
Here, $\hat{\chi}(x)$, $\hat{R}(x)$, and $\hat{U}(x)$ are called the scaling functions, and $\nu$, $\gamma$ and $\beta$ are the critical exponents $\xi\sim|\tilde{J}-\tilde{J}_{c}|^{-\nu}$ where $\xi$ is the correlation length, $\chi\sim|\tilde{J}-\tilde{J}_{c}|^{-\gamma}$, and $R\sim|\tilde{J}-\tilde{J}_{c}|^{\beta}$ (for $\tilde{J}>\tilde{J}_{c}$). The scaling functions are independent of $L$ and converge to a finite value at the critical point (i.e. $\hat{R}(0)$, $\hat{\chi}(0)$ and $\hat{U}(0)$ are constants). In what follows we assume that the finite-size scaling theory holds out of equilibrium, as is often done [189]. We set $\tilde{K}=0.1$ where numerical evidence shows that the transition is continuous (like a second-order transition).
@@
The critical point $\tilde{J}_{c}$ is found using the Binder cumulant. Away from the critical point, the value of $U$ depends on system size, approaching $1/3$ in the disordered and $2/3$ in the ordered regime as $L\to\infty$ [190, 191]. However, at the critical point $U$ approaches a fixed value $U(\tilde{J}_{c})=U^{*}$ independent of $L$. Hence, the critical point is found from the intersection point of all the $U$ vs. $\tilde{J}$ curves plotted for different system sizes. Figure 14a shows that indeed all the curves cross at a single point ($\tilde{J}_{c}=1.34614\pm 0.00001$ for $\tilde{K}=0.1$).
@@
In addition to $U$, we plot $\chi$ and $R$ as a function of $\tilde{J}$, see Fig. 14a-c. The data points are interpolated using a cubic spline method with ‘not-a-knot’ boundary conditions as implemented in scipy.interpolate.CubicSpline [192]. For each linear system size $L$, we find the slope of $U$ at the critical point, ${\rm d}U/{\rm d}\tilde{J}|_{\tilde{J}=\tilde{J}_{c}}(L)$, the maximum of $\chi$, $\chi_{\rm max}(L)$, and the value of $R$ at the critical point, $R(\tilde{J}=\tilde{J}_{c},L)$. We then use the finite-size scaling relations (which can be derived from Eq. (57[186, 187])
$\begin{split}\frac{{\rm d}U}{{\rm d}\tilde{J}}|_{\tilde{J}=\tilde{J}_{c}}(L)&=aL^{1/\nu}\\ \chi_{\rm max}(L)&=bL^{\gamma/\nu}\\ R(\tilde{J}=\tilde{J}_{c},L)&=cL^{-\beta/\nu}\end{split}$ (58)
@@
and perform linear fits by taking the logarithm of both sides of the equations (Fig. 14(d-f)). The procedure outlined here yields the critical exponents. To ensure good accuracy, all averages are taken using $10^{6}$ samples. To make the samples approximately independent, they are separated by two correlation times of $s$ (the correlation times were computed at $\tilde{J}=1.346$). Simulation runs are initialized in a fully disordered state. The error is estimated using the bootstrap method [186, 193]: the same procedure is repeated for multiple samples where each sample is constructed by drawing $10^{6}$ measurements of $s$ randomly from our data while allowing duplicates. The standard deviation of the critical exponents is estimated by the standard deviation calculated from $10^{5}$ bootstrap resamplings of the data (Fig. 14(g-i)). This estimated error is due to finite sampling but does not include other systematic errors (e.g. finite-size corrections to Eqs. (5758[187]). Our results are
$\begin{split}\nu&=0.675\pm 0.005\\ \gamma&=1.328\pm 0.009\\ \beta&=0.347\pm 0.002\,,\\ \end{split}$ (59)
which are all in good agreement with 3D XY critical exponents $\nu_{XY}=0.672$, $\gamma_{XY}=1.318$, $\beta_{XY}=0.349$ [194, 195], more so than with 3D Ising critical exponents $\nu_{I}=0.630$, $\gamma_{I}=1.237$, $\beta_{I}=0.326$ [194], see graphical comparison in Fig. 15. Further evidence that the disorder-to-swap phase transition belongs to the same universality class as the disorder-to-order XY phase transition is given by the value of the Binder cumulant as the critical point, $U({\tilde{J}_{c}})=U^{*}$, which is also universal [188]. We find
$U^{*}=0.5877\pm 0.0002\,,$ (60)
consistent with studies on the 3D XY model reporting $U^{*}\approx 0.586$ [196]. In contrast, the 3D Ising model yields $U^{*}\approx 0.467$ [197].
@@
Fig. 14: Calculation of the critical exponents of the 3D NR Ising model. We set $\tilde{K}=0.1$. (a-c) The Binder cumulant $U$, the susceptibility $\chi$, and the order parameter $R$ as a function of $\tilde{J}$ for linear system sizes $L=20,24,28,32,36,40,44,48$. The vertical dashed grey line indicates the critical point, $\tilde{J}_{c}=1.34614$ while the horizontal dotted black line indicates that value of the Binder cumulant at the critical point, $U^{*}=0.5877$. Each point is calculated by averaging over $10^{6}$ samples of $s$. (d-f) ${\rm d}U/{\rm d}{\tilde{J}}|_{\tilde{J}=\tilde{J}_{c}}$, $\chi_{\rm max}$ and $R(\tilde{J}_{c})$ as a function of $L$ on a log-log scale. Dots are extracted from analyzing panels a-c and full lines indicate best linear fits. (g-i) Standard error of $\nu$, $\gamma$, and $\beta$ as a function of the number of bootstrap samples ($10^{5}$ bootstrap resamplings are performed in total).
@@
Fig. 15: Compatibility with XY universality class. The critical exponents and the Binder cumulant at the critical point of the 3D nonreciprocal Ising model (red) for $\tilde{K}=0.1$ extracted from finite-size scaling analysis along with their standard deviation represented by a semi-transparent red rectangle, and compared with the 3D Ising model (blue) and the 3D XY model (green).